a) We are planning to operate a mixed flow reactor to convert A into R. This is a liquid phase reaction, with the stoichiometryA→R. The rate of reaction is given below. What size of mixed flow reactor is needed to achieve 75% conversion of a feed stream of 1000 moles of A per hour?

CA,mol/l. 0.1 0.2 0.3 04 0.5 0.6 0.7 0.8 1 1.3 2. -rA,mol/(l.min). 0.1 0.3 0.5 0.6 0.5 0.25 0.1 0.06 0.05 0.045 0.042

Solve this problem with step by step process?


1
Expert's answer
2021-10-11T07:01:18-0400


Size of mixed flow reactor OF 1000 moles of A per hour:

We can compute this as follows:We obtain:

0.469dm3/mol


"C_A=C_{A0}(1-X)"


"C_B=C_{A0}\frac{1}{2} X"


"K_{Ce}=\frac{X_e}{2C_{A0}(1-X_e)^2}"


"=\frac{0.6}{8(0.4)^2}"


"=0.469 dm^3\/mol"


We obtain:

0.469dm3/mol

It means we require mixed flow reactor OF 1000 moles of A to be 0.469dm3/mol every hour.